How would I simplify this problem.
5 - 2{5-2(2x+4)}

Answer by tutorcecilia(2152)

5 - 2{5-2(2x+4)} [follow the order of operations (PEMDAS)]
5 - 2{5-4x-8)
5 - 2{-3-4x)
5 +6+8x
11+8x
11+8x=0 [Set the expression equal to zero)
11-11+8x=0-11
8x=-11
8x/8=-11/8
x=-11/8
.
Check by plugging (x-11/8) back into the original expression
5 - 2{5-2(2x+4)}=0
5 - 2{5-2(2(-11/8)+4)}=0
0=0